• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mitra

    Log entry time 08:07:18 on May 7,2011

    Entry number 358455

    keyword=Owl shift summary

    Shift crew:
    -------------
    SL: Mitra Shabestari
    TO: Chao GU
    3rd: Zinzhan Bai


    00:00 --- Change to kinematics 3.1 in progress.
    00:20 --- Moved target to BeO
    00:50 --- Moved target to Optics
    01:15 --- Moved target to Ca40
    01:40 --- Lost NMR lock on left HRS
    02:19 --- Moved target to Ca48
    02:20 --- Lost NMR lock on left HRS
    03:23 --- I noticed that BPMA fluctuates, so asked MCC to set the X and Y values.
    03:30 --- Started changing kinematics to 6.5. Moved target to He3.
    04:10 --- Started kinematics 6.5
    05:04 --- During run 4251 I noticed that there is no start of the run. ROC8 8 was frozen.
    After rebooting ROC8, ROC1 was not responding. Rebooting ROC1 didn't help, so
    we had to reboot CODA a few times to get it running.
    05:37 --- Smooth production run till the end of the shift


    Notes:
    Left arm dipole lost its NMR lock several times during kinematics 3.1;
    it takes 2-5 minutes to lock the NMR.


    Run list for kinematics 3.1
    =============================

    Run# --- target, current, #of events

    4242 --- BeO, 5muA, 54K
    4243 --- BeO, 5muA, 123K
    4244 --- Optics, 5muA, 132K
    4245 --- Ca40, 40muA, 4.4M
    4246 --- Ca40, 40muA, 3.2M
    4247 --- Ca48, 40muA, 4.0M
    4248 --- Ca48, 40muA, 4.0M
    4249 --- Ca48, 40muA, 4.1M
    4250 --- Ca48, 40muA, 5.0M


    Run list for kinematics 5.6
    =============================

    Run# --- target, current, #of events

    4251 --- He3, 120muA, No start of the run. ROC8 was frozen.
    4252 --- He3, 120muA, 1.0M
    4253 --- He3, 120muA, 912K
    4254 --- He3, 120muA, ... (ongoing)